# 5- V o l u m e 1, N u m b e r 1, S e p t e m b e lr 2 0 1 8

### v2

@@
[…]
5- V o l u m e 1, N u m b e r 1, S e p t e m b e lr 2 0 1 8
[…]
3 DoF Analysis and Beamforming Design for Active IRS-aided Multi-user MIMO Wireless Communication in Rank-deficient Channels
@@
Feng Shu, Jinbing Jiang, Xuehui Wang*, Ke Yang, Chong Shen
Qi Zhang, Dongming Wang and Jiangzhou Wang
[…]
Key words:
DoF, IRS, sum-rate, MU-MIMO, beamforming
[…]
6
@@ 1 Introduction
Fortunately, as the improvement of meta-materials, a novel green forwarder called passive intelligent reflecting surface (IRS) is recently proposed [3]. Its surface integrates a multitude of low-cost passive reflective elements, by adjusting the phase shift of each element, IRS can intelligently control the wavefront of the incident signals in a desired manner. In terms of hardware cost, rate gain and channel reconfigurability, IRS has many advantages such as low circuit cost, high energy efficiency, rate enhancement, and coverage extension. Thus, it is viewed as one of the most potential key techniques for the future wireless networks like 6G, UAV communications, satellite communications, and marine communications. Because of the unique features, IRS can be applied to wireless networks for coverage extension and energy conservation [4]-[5]. Besides that, the hot noise introduced in IRS can be negligible [6]. For these reasons, more and more researchers pay attention to IRS [7]-[8]. Many authors used the advantage of IRS to control the direction of incident signals to assist communication networks for secure transmission [9]-[10]. In an IRS-assisted directional modulation (DM) network, the authors in [11] designed the beamforming to enhance the secrecy rate performance. To enhance the energy harvesting performance (EHP), an IRS-aided simultaneous wireless information and power transfer (SWIPT) wireless network was put forward in [12], while it was demonstrated that the EHP of the SWIPT system could be enhanced by IRS. In [13], IRS was adopted to mobile edge computing system to reduce the computational latency. The authors employed IRS to cognitive radio systems in [14], where the problem of maximizing the weighted sum rate (WSR) of secondary users (SUs) was formulated. By jointly optimizing SU transmit precoding and IRS phase shifts, it was showed that the significant WSR improvement could be attained with the aid of IRS. Moreover, by controlling the incident light beam, IRS could extend the range of rotation angle of the visible light communication receiver [15], while the field-of-view was improved. An IRS-aided decode-and-forward (DF) relay network was designed in [16], where a alternately iterative structure was proposed to obtain the analysis solutions to DF beamforming vector and IRS reflection matrix, so that the receive power was maximum.
@@
Based on these unique advantages, a lot of research on active IRS has been implemented [22]-[24]. The authors in [25] compared the active and passive IRS-aided system theoretically and numerically, and the simulation results showed that active IRS was better than passive when the power budget was not so small and the number of IRS elements was not so large. In [26], the authors have demonstrated that introducing active IRS to wireless network can obtain rate improvement compared to passive IRS-aided network. An active IRS-aided multi-user multiple-input multiple-output (MU-MIMO) was presented in [27], where energy efficiency was maximized by jointly designing transmit beamforming and energy-included IRS reflection matrix. The authors in [28] investigated a new active IRS-aided uplink multi-antenna non-orthgonal multiple access system, base on minimum mean square error (MMSE) criterion, the proposed system achieved higher sum rate than those of various benchmarks by an efficient alternating optimization (AO) algorithm. With the practical hardware impairments, [29] considered an active IRS-assisted the multiuser downlink transmission systems. By designing BS transmit beamforming and active IRS reflection matrix including the amplification factors and the phase shifts, the maximum sum date rate could be achieved. To improve the rate of a wireless network aided by an active IRS, [30] first simplified signal-to-noise ratio (SNR), and applied Rayleigh-Ritz theorem to maximize SNR by designing active IRS reflection matrix. To improve rate, generalized maximum ratio reflection method and maximizing SNR method based on fractional programming were respectively proposed.
@@
    To evaluate the DoF gains achieved by active IRS, an IRS-aided multi-user network is established and the corresponding achievable DoF upper bounds in single-user MIMO network is derived in rank-deficient channels, respectively. Then, the corresponding results are also extended to the multi-user scenarios. In particular, it is found that introducing IRS will make a dramatic DoF enhancement. For example, in a general channel, the number of DoFs of IRS-aided wireless network is twice that of no IRS. However, for LoS channel, as a low-rank channel, the corresponding DoF gain will become trivial as the number $K$ of users tends to medium or large-scale. When the $K$ goes to 1, the DoF enhancement is close or equal to that of the rank-deficient channel. Finally, the special condition for achieving the maximum DoF of single-user in LoS channel is constructed, i.e. channels orthogonal condition (OC) to achieve two DoFs .
@@
    To confirm the sum-rate gains achieved by incremental DoF created by IRS in rank-deficient channel, a two-layer leakage plus minimum mean square error (TLL-MMSE) method is proposed, where the precoding vectors and receive beamforming vectors are designed by the leakage concept and MMSE rule, respectively, while the phase is adjusted by the leakage principle. Simulation results shows that the sum-rate of the proposed TLL-MMSE with IRS may make a significant sum-rate improvement over no IRS, i. e., achieving up to more than 100% rate gain over no IRS.
@@ 2 System Model
$\mathbf{x}=\sum_{i=1}^{K}\mathbf{V}_{i}\mathbf{s}_{i},$ (1)
where $\mathbf{V}_{i}\in\mathbb{C}^{M\times L_{i}}$ is the $i$-th transmission beamforming matrix at BS, and $\mathbf{s}_{i}\in\mathbb{C}^{L_{i}\times 1}$ is the corresponding data symbol, where $\mathbb{E}\{\mathbf{s}_{i}^{H}\mathbf{s}_{i}\}=1$. The reflect signal of the IRS is
$\mathbf{y}^{t}=\mathbf{\Theta}_{k}\mathbf{H}_{BI}\mathbf{x}+\mathbf{\Theta}_{k}\mathbf{n},$ (2)
where $\mathbf{H}_{BI}\in\mathbb{C}^{N\times M}$ denotes the channel matrix from BS to IRS, and $\mathbf{\Theta}_{k}=\mathrm{diag}(\alpha_{1}e^{j\theta_{1}},...,\alpha_{N}e^{j\theta_{N}})\in\mathbb{C}^{N\times N}$ stands for the reflection coefficient matrix of IRS, where $\alpha_{l}$ represents amplitude gain and $\theta_{l}$ stands for phase shift, respectively. And $\mathbf{n}\in\mathbb{C}^{N\times 1}$ is the additive white Gaussian noise (AWGN) at IRS , and $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}\mathbf{I}_{N})$.
@@
$\mathbf{y}_{k}^{r}=$ $\mathbf{H}_{dk}\sum_{i=1}^{K}\mathbf{V}_{i}\mathbf{s}_{i}+\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\sum_{i=1}^{K}\mathbf{V}_{i}\mathbf{s}_{i}$
$+\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{n}+\mathbf{z}_{k},$ (3)
where $\mathbf{z}_{k}\sim\mathcal{CN}(0,\sigma_{z}^{2}\mathbf{I}_{N})$ After the receive beamforming is performed, we have
$\mathbf{y}_{k}$ $=\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\mathbf{V}_{k}\mathbf{s}_{k}+\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\mathbf{V}_{k}\mathbf{s}_{k}$
$+\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\sum_{i=1,i\neq k}^{K}\mathbf{V}_{i}\mathbf{s}_{i}+\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{n}$
$+\mathbf{U}_{k}^{H}\mathbf{z}_{k}+\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\sum_{i=1,i\neq k}^{K}\mathbf{V}_{i}\mathbf{s}_{i},$ (4)
where $\mathbf{U}_{k}^{H}\in\mathbb{C}^{L_{k}\times Q_{k}}$ is the receive beamforming matrix , $\mathbf{G}_{k}^{H}\in\mathbb{C}^{Q_{k}\times N}$ and $\mathbf{H}_{dk}\in\mathbb{C}^{Q_{k}\times M}$ denote the channel gain matrices from IRS to user $k$ and BS to user $k$, respectively.
@@
$\mathbb{E}\{\mathbf{y}_{k}^{H}\mathbf{y}_{k}\}=\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\mathbf{V}_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}\mathbf{V}_{k}^{H}\mathbf{H}_{dk}^{H}\mathbf{U}_{k}$
$+\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\mathbf{V}_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}\mathbf{V}_{k}^{H}\mathbf{H}_{BI}^{H}\mathbf{\Theta}_{k}^{H}\mathbf{G}_{k}\mathbf{U}_{k}$
$+\sum_{i=1,i\neq k}^{K}\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\mathbf{V}_{i}\mathbf{s}_{i}\mathbf{s}_{i}^{H}\mathbf{V}_{i}^{H}\mathbf{H}_{dk}^{H}\mathbf{U}_{k}$
$+\sum_{i=1,i\neq k}^{K}\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\mathbf{V}_{i}\mathbf{s}_{i}\mathbf{s}_{i}^{H}\mathbf{V}_{i}^{H}\mathbf{H}_{BI}^{H}\mathbf{\Theta}_{k}^{H}\mathbf{G}_{k}\mathbf{U}_{k}$
$+\sigma^{2}\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{\Theta}_{k}^{H}\mathbf{G_{k}}\mathbf{U}_{k}+\sigma^{2}_{z}\mathbf{U}_{k}^{H}\mathbf{U}_{k}.$ (5)
Then the rate of user $k$ can be obtained
$\mathbf{\gamma}_{k}$ $=\log_{2}\left|\mathbf{I}_{k}+\frac{\mathbf{A}_{1}+\mathbf{A}_{2}}{\mathbf{B}_{1}+\mathbf{B}_{2}+\mathbf{B}_{3}}\right|,$ (6)
@@
$\mathbf{A}_{2}$ $=\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\mathbf{C}\mathbf{H}_{BI}^{H}\mathbf{\Theta}_{k}^{H}\mathbf{G}_{k}\mathbf{U}_{k},$
@@
$\mathbf{B}_{2}$ $=\sum_{i=1,i\neq k}^{K}\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{H}_{BI}\mathbf{C}\mathbf{H}_{BI}^{H}\mathbf{\Theta}_{k}^{H}\mathbf{G}_{k}\mathbf{U}_{k},$
$\mathbf{B}_{3}$ $=\sigma^{2}\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}_{k}\mathbf{\Theta}_{k}^{H}\mathbf{G_{k}}\mathbf{U}_{k}+\sigma^{2}_{z}\mathbf{U}_{k}^{H}\mathbf{U}_{k},$
$\mathbf{C}$ $=\mathbf{V}_{k}\mathbf{s}_{k}\mathbf{s}_{k}^{H}\mathbf{V}_{k}^{H}.$ (7)

### v3

# TSINGHUA  SCIENCE  AND  TECHNOLOGY

@@
Abstract: Due to its ability of significantly improving data rate, intelligent reflecting surface (IRS) will be a potential crucial technique for the future generation wireless networks like 6G. In this paper, we will focus on the analysis of degree of freedom (DoF) in IRS-aided multi-user MIMO network. Firstly, the DoF upper bound of IRS-aided single-user MIMO network, i.e., the achievable maximum DoF of such a system, is derived, and the corresponding results are extended to the case of IRS-aided multiuser MIMO by using the matrix rank inequalities. In particular, in serious rank-deficient, also called low-rank, like line-of-sight (LoS) channel, the network DoF may doubles over no-IRS with the help of IRS. To verify the rate performance gain from augmented DoF, three closed-form beamforming methods, null-space projection plus maximize transmit power and maximize receive power (NSP-MTP-MRP), Schmidt orthogonalization plus minimum mean square error (SO-MMSE) and two-layer leakage plus MMSE (TLL-MMSE) are proposed to achieve the maximum DoF. Simulation results shows that IRS does make a dramatic rate enhancement. For example, in a serious rank-deficient channel, also called low-rank, the sum-rate of the proposed TLL-MMSE aided by IRS is up to 2.54 times that of no IRS. This means that IRS may make a significant DoF improvement in such a channel.
Key words:
DoF, IRS, sum-rate, MU-MIMO, beamforming
$\bullet$ Jinbing Jiang, Ke Yang, Chong Shen and Qi Zhang are with the School of Information and Communication Engineering, Hainan University, Haikou, 570228, China. Email:JJB2511@163.com;;23220854100060@hainanu.edu.cn;chongshen@haina
    nu.edu.cn;hdzhangqi0509@163.com.
$\bullet$ F. Shu is with the School of Information and Communication Engineering, Hainan University, Haikou, 570228, China, and also with the School of Electronic and Optical Engineering, Nanjing University of Science and Technology, Nanjing, 210094, China. Email: shufeng0101@163.com.
$\bullet$ Xuehui Wang is with the School of Mathematics and Statistics, Hainan Normal University, Haikou, 571127, China. E-mail: wangxuehui0503@163.com.
$\bullet$ Dongming Wang is with the National Mobile Communications Research Laboratory, Southeast University, Nanjing 210096, China. E-mail: wangdm@seu.edu.cn.
$\bullet$ Jiangzhou Wang is with the School of Engineering, University of Kent, Canterbury CT2 7NT, U.K., Email: j.z.wang@kent.ac.uk.
@@
Manuscript received: 15-Oct-2024; revised: 26-Mar-2025; accepted: 22-Apr-2025
@@ 1 Introduction
Fortunately, as the improvement of meta-materials, a novel green forwarder called passive intelligent reflecting surface (IRS) is recently proposed [3]. Its surface integrates a multitude of low-cost passive reflective elements, by adjusting the phase shift of each element, IRS can intelligently control the wavefront of the incident signals in a desired manner. In terms of hardware cost, rate gain and channel reconfigurability, IRS has many advantages such as low circuit cost, high energy efficiency, rate enhancement, and coverage extension. Thus, it is viewed as one of the most potential key techniques for the future wireless networks like 6G, UAV communications, satellite communications, and marine communications. Because of the unique features, IRS can be applied to wireless networks for coverage extension and energy conservation [4]-[5]. Besides that, the hot noise introduced in IRS can be negligible [6]. For these reasons, more and more researchers pay attention to IRS [7]-[8]. Many authors used the advantage of IRS to control the direction of incident signals to assist communication networks for secure transmission [9]-[10]. In an IRS-assisted directional modulation (DM) network, the authors in [11] designed the beamforming to enhance the secrecy rate performance. To enhance the energy harvesting performance (EHP), an IRS-aided simultaneous wireless information and power transfer (SWIPT) wireless network was put forward in [12], while it was demonstrated that the EHP of the SWIPT system could be enhanced by IRS. In [13], IRS was adopted to mobile edge computing system to reduce the computational latency. The authors employed IRS to cognitive radio systems in [14], where the problem of maximizing the weighted sum-rate (WSR) of secondary users (SUs) was formulated. By jointly optimizing SU transmit precoding and IRS phase shifts, it was showed that the significant WSR improvement could be attained with the aid of IRS. Moreover, by controlling the incident light beam, IRS could extend the range of rotation angle of the visible light communication receiver [15], while the field-of-view was improved. An IRS-aided decode-and-forward (DF) relay network was designed in [16], where a alternately iterative structure was proposed to obtain the analysis solutions to DF beamforming vector and IRS reflection matrix, so that the receive power was maximum.
@@
Based on these unique advantages, a lot of research on active IRS has been implemented [22]-[24]. The authors in [25] compared the active and passive IRS-aided system theoretically and numerically, and the simulation results showed that active IRS was better than passive when the power budget was not so small and the number of IRS elements was not so large. In [26], the authors have demonstrated that introducing active IRS to wireless network can obtain rate improvement compared to passive IRS-aided network. An active IRS-aided multi-user multiple-input multiple-output (MU-MIMO) was presented in [27], where energy efficiency was maximized by jointly designing transmit beamforming and energy-included IRS reflection matrix. The authors in [28] investigated a new active IRS-aided uplink multi-antenna non-orthgonal multiple access system, base on minimum mean square error (MMSE) criterion, the proposed system achieved higher sum-rate than those of various benchmarks by an efficient alternating optimization (AO) algorithm. With the practical hardware impairments, [29] considered an active IRS-assisted the multiuser downlink transmission systems. By designing BS transmit beamforming and active IRS reflection matrix including the amplification factors and the phase shifts, the maximum sum date rate could be achieved. To improve the rate of a wireless network aided by an active IRS, [30] first simplified signal-to-noise ratio (SNR), and applied Rayleigh-Ritz theorem to maximize SNR by designing active IRS reflection matrix. To improve rate, generalized maximum ratio reflection method and maximizing SNR method based on fractional programming were respectively proposed. The authors in [31] proposed a multi-IRS-aided multi-stream DM network. To eliminate the interference among different data streams, the null-space projection (NSP) at BS and receive zero-forcing (ZF) at user is adopted to design the beamforming of transmitter and receiver, and phase alignment (PA) are adopted to obtain the phase shift matrix (PSM). In [32], an active IRS-assisted multi-user system was considered, and a method of minimizing the BS transmit power was proposed to jointly optimizes the reflection matrix at IRS and the beamforming vectors at BS to achieve a green wireless communication. [33] proposed a low-complexity adaptive selection beamforming scheme for the practical implementation of IRS-assisted single-user wireless networks. Compared with AO, the proposed scheme requires lower computational cost.
@@
    To evaluate the DoF gains achieved by active IRS, an IRS-aided multi-user network is established and the corresponding achievable DoF upper bounds in single-user MIMO network is derived in rank-deficient channels. Then, the corresponding results are also extended to the multi-user scenarios. In particular, it is found that introducing IRS will make a dramatic DoF enhancement. For example, in a general channel, the number of DoFs of IRS-aided wireless network is twice that of no IRS. However, for LoS channel, as a low-rank channel, the corresponding DoF gain will become trivial as the number $K$ of users tends to medium or large-scale. When the $K$ goes to 1, the DoF enhancement is close or equal to that of the rank-deficient channel. Finally, the special condition for achieving the maximum DoF of single-user in LoS channel is constructed, i.e. channels orthogonal condition (OC) to achieve two DoFs .
@@
    To confirm the sum-rate gains achieved by incremental DoF created by IRS in rank-deficient channel, a two-layer leakage plus minimum mean square error (TLL-MMSE) method is proposed, where the precoding vectors and receive beamforming vectors are designed by the leakage concept and MMSE rule, respectively, while the phase is adjusted by the leakage principle. Simulation results shows that the sum-rate of the proposed TLL-MMSE with IRS may make a significant sum-rate improvement over no IRS, i. e., achieving up to 2.54 times gain over no IRS.
@@ 2 System Model
$\mathbf{x}=\sum_{k=1}^{K}\mathbf{V}_{k}\mathbf{s}_{k},$ (1)
where $\mathbf{V}_{k}\in\mathbb{C}^{M\times L_{k}}$ is the $k$-th transmission beamforming matrix at BS, and $\mathbf{s}_{k}\in\mathbb{C}^{L_{k}\times 1}$ is the corresponding data symbol, where $\mathbb{E}\{\mathbf{s}_{k}^{H}\mathbf{s}_{k}\}=\mathbf{I}_{L_{k}}$. The reflect signal of the IRS is
$\mathbf{y}^{t}=\mathbf{\Theta}\mathbf{H}_{BI}\mathbf{x}+\mathbf{\Theta}\mathbf{n},$ (2)
where $\mathbf{H}_{BI}\in\mathbb{C}^{N\times M}$ denotes the channel matrix from BS to IRS, and $\mathbf{\Theta}=\mathrm{diag}(\alpha_{1}e^{j\theta_{1}},...,\alpha_{N}e^{j\theta_{N}})\in\mathbb{C}^{N\times N}$ stands for the reflection coefficient matrix of IRS, where $\alpha_{l}$ represents amplitude gain and $\theta_{l}$ stands for phase shift, respectively. And $\mathbf{n}\in\mathbb{C}^{N\times 1}$ is the additive white Gaussian noise (AWGN) at IRS, and $\mathbf{n}\sim\mathcal{CN}(0,\sigma^{2}_{irs}\mathbf{I}_{N})$.
@@
$\mathbf{y}_{k}^{r}=$ $\mathbf{H}_{dk}\sum_{i=1}^{K}\mathbf{V}_{i}\mathbf{s}_{i}+\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\sum_{i=1}^{K}\mathbf{V}_{i}\mathbf{s}_{i}$
$+\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{n}+\mathbf{z}_{k},$ (3)
where $\mathbf{G}_{k}^{H}\in\mathbb{C}^{Q_{k}\times N}$ and $\mathbf{H}_{dk}\in\mathbb{C}^{Q_{k}\times M}$ denote the channel gain matrices from IRS to user $k$ and BS to user $k$, respectively, $\mathbf{z}_{k}\in\mathbb{C}^{Q_{k}\times 1}$ is the additive white Gaussian noise (AWGN) at user $k$, and $\mathbf{z}_{k}\sim\mathcal{CN}(0,\sigma_{z}^{2}\mathbf{I}_{N})$. After the receive beamforming is performed, we have
$\mathbf{y}_{k}$ $=\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\mathbf{V}_{k}\mathbf{s}_{k}+\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\mathbf{V}_{k}\mathbf{s}_{k}$
$+\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\sum_{i=1,i\neq k}^{K}\mathbf{V}_{i}\mathbf{s}_{i}+\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{n}$
$+\mathbf{U}_{k}^{H}\mathbf{z}_{k}+\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\sum_{i=1,i\neq k}^{K}\mathbf{V}_{i}\mathbf{s}_{i},$ (4)
where $\mathbf{U}_{k}^{H}\in\mathbb{C}^{L_{k}\times Q_{k}}$ is the receive beamforming matrix , .
@@
$\mathbb{E}\{\mathbf{y}_{k}^{H}\mathbf{y}_{k}\}=\|\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\mathbf{V}_{k}\mathbf{V}_{k}^{H}\mathbf{H}_{dk}^{H}\mathbf{U}_{k}\|_{F}$
$+\|\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\mathbf{V}_{k}\mathbf{V}_{k}^{H}\mathbf{H}_{BI}^{H}\mathbf{\Theta}^{H}\mathbf{G}_{k}\mathbf{U}_{k}\|_{F}$
$+\sum_{i=1,i\neq k}^{K}\|\mathbf{U}_{k}^{H}\mathbf{H}_{dk}\mathbf{V}_{i}\mathbf{V}_{i}^{H}\mathbf{H}_{dk}^{H}\mathbf{U}_{k}\|_{F}$
$+\sum_{i=1,i\neq k}^{K}\|\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{H}_{BI}\mathbf{V}_{i}\mathbf{V}_{i}^{H}\mathbf{H}_{BI}^{H}\mathbf{\Theta}^{H}\mathbf{G}_{k}\mathbf{U}_{k}\|_{F}$
$+\sigma_{irs}^{2}\|\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}\mathbf{\Theta}\mathbf{\Theta}^{H}\mathbf{G_{k}}\mathbf{U}_{k}\|_{F}+\sigma^{2}_{z}\|\mathbf{U}_{k}^{H}\mathbf{U}_{k}\|_{F}.$ (5)
Then the corresponding rate can be obtained as
$\mathbf{\gamma}_{k}$ $=\log_{2}\left|\mathbf{I}_{k}+(\mathbf{A}_{1}+\mathbf{A}_{2})(\mathbf{B}_{1}+\mathbf{B}_{2}+\mathbf{B}_{3})^{-1}\right|,$ (6)
@@
$\mathbf{A}_{2}$ $=\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}{\mathbf{\Theta}}\mathbf{H}_{BI}\mathbf{C}\mathbf{H}_{BI}^{H}{\mathbf{\Theta}}^{H}\mathbf{G}_{k}\mathbf{U}_{k},$
@@
$\mathbf{B}_{2}$ $=\sum_{i=1,i\neq k}^{K}\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}{\mathbf{\Theta}}\mathbf{H}_{BI}\mathbf{C}\mathbf{H}_{BI}^{H}{\mathbf{\Theta}}^{H}\mathbf{G}_{k}\mathbf{U}_{k},$
$\mathbf{B}_{3}$ $={\sigma^{2}_{irs}}\mathbf{U}_{k}^{H}\mathbf{G}_{k}^{H}{\mathbf{\Theta}}{\mathbf{\Theta}}^{H}\mathbf{G_{k}}\mathbf{U}_{k}+\sigma^{2}_{z}\mathbf{U}_{k}^{H}\mathbf{U}_{k},$
$\mathbf{C}$ $=\mathbf{V}_{k}\mathbf{V}_{k}^{H}.$ (7)